Ivonne has bought shares of RIO, Inc. stock for $25.00 per share. She expects a 1.00 dividend at the end of this year. After 2 years, she expects to receive a dividend of $1.25 and to sell the stock for $28.75. What is Ivonne's required rate of return?

  • 11. 6%

  • 12.66

  • 13.13%

  • Markhem Enterprises is expected to earn $1.34 per share this year. The company has a dividend payout ratio of 40% and a P/E ratio of 18. What should one share of common stock in Markhem Enterprises be selling for in the market?

  • $24.12

  • $33

  • The risk-free rate of return is 4.2 percent, the expected market return is 9 percent, and the beta for Lea, Inc. is 1.12. What is Lea's required rate of return?

  • 9.58%

  • 13%
Answer #1

1. Ivonne's gain is the dividends and the capital gains received after selling the share.

Cash outflow in Year 0 = -25

Cash inflow in year 1 = +1

Cash inflow in year 2 = 1.25 + 28.75 = 30

IRR = 11.56 = 11.6%

2. P/E ratio = Price / Earnings

18 = Price/ 1.34

Price = $24.12

3. CAPM = Rf + b[ E(m) - Rf]

= 4.2 +1.12 ( 9 - 4.2 )

= 9.58%